# Multiple ergodic averages along functions from a Hardy field: convergence, recurrence and combinatorial applications

### v2

@@ 5.2.   Dealing with rational polynomials
5.3 holds if we further assume that each function $f_{i}$ satisfies the condition
If |$\text{If }|f_{i}(t)-p(t)|\to 0\text{ for some }p\in\mathbb{R}[t]\text{ with }p-p(0)\in\mathbb{Q}[t]\text{, then }p-p(0)\in\mathbb{Z}[t].$ (5.1)
To reduce 5.3 to 5.4 we need the following lemma.
@@
Because $W\in\mathcal{H}$, $\lim_{t\to\infty}W(t)/t$ exists and, because $W(t)\ll t$, that limit is in $[0,\infty)$. If it is nonzero, then $L:=\lim\frac{W(Rt+d)}{W(t)}=\lim\frac{W(Rt+d)}{Rt+d}\frac{t}{W(t)}\frac{Rt+d}{t}=R$, and if $\lim W(t)/t=0$ then $W(t)/t$ is eventually decreasing, so $L\leqslant R$. Since $W$ is eventually increasing (because $1\prec W(t)$) we have $L\geqslant 1$. Next it follows from L’Hôpital’s rule and the mean value theorem that $w(Rt+d)/w(t)\to\frac{L}{R}$.
@@
Observe that the polynomials $p_{i,d}\colon t\mapsto p_{i}(Rt+d)-p_{i}(d)$ have integer coefficients for any fixed $d\in\mathbb{N}$ and any $i\leqslant m$. Therefore, for each $d\in\mathbb{N}$ and $i\in\{1,\dots,k\}$ the function $f_{i,d}\colon t\mapsto f_{i}(Rt+d)$ satisfies condition (5.1). Parts 1 and 2 of 5.3 now follow from 5.4 and 5.5.
Finally, we prove part 3. Suppose that the functions $f_{1},\ldots,f_{k}$ satisfy condition 2 of A and let $q_{1},\dots,q_{\ell}\in\mathbb{Z}[t]$ be jointly intersective and such that $\mathrm{poly}(f_{1},\dots,f_{k})\subset\mathrm{span}^{*}(q_{1},\dots,q_{\ell})$. By the pigeonhole principle, for some $d\in\{0,\dots,R-1\}$ the polynomials $\tilde{q}_{i}:t\mapsto q_{i}(Rt+d)$, $i=1,\dots,\ell$ are jointly intersective. Therefore the functions $f_{1,d},\ldots,f_{k,d}$ also satisfy condition 2 of A. Then using 5.5 and 5.4 we have
@@ 5.3.   Reduction to a statement about uniform distribution
Thus far, we have reduced 5.1 to 5.4. Our next goal is to reduce 5.4 to the following result:
@@
Let $G$ be a connected and simply connected nilpotent Lie group, let $\Gamma\subset G$ be a uniform and discrete subgroup, let $X=G/\Gamma$, and let $a\in G$. Let $f_{1},\dots,f_{k}\in\mathcal{H}$ satisfying property 1 for some $W\in\mathcal{H}$ with $1\prec W(t)\ll t$ and property (5.1) from 5.4. Then there exists a Borel probability measure $\nu$ on $X^{k}$ such that the sequence
@@
Moreover, if condition 2 of A is satisfied then the point $1_{X^{k}}=1_{G^{k}}\Gamma^{k}$ belongs to the support of $\nu$, and if condition 1 of A is satisfied then $\nu$ is the Haar measure on the subnilmanifold $Y^{k}\subset X^{k}$, where $Y=\overline{\{a^{n}\Gamma:n\in\mathbb{Z}\}}$.
Proof of 5.4 assuming 5.6.
Let $G$ be a connected and simply connected nilpotent Lie group and let $\Gamma\subset G$ be a uniform and discrete subgroup such that $X=G/\Gamma$ and $T:X\to X$ is given by left multiplication by some element $a\in G$. Our goal is to show that
  1. [label=(),ref=(),leftmargin=*]
  2. 1.
@@
  3. 2.
    If $f_{1},\ldots,f_{k}$ satisfy condition 1 of A then for any $h_{1},\ldots,h_{k}\in C(X)$ and any $x\in X$
@@
  4. 3.
    If $f_{1},\ldots,f_{k}$ satisfy condition 2 of A then for any $A\in\mathcal{B}$ with $\mu(A)>0$ we have
[…]
@@
Now let $\nu_{x}$ be the measure on $X^{k}$ given by 5.6 with $g_{x}^{-1}ag_{x}$ in the role of $a$ and define $H_{x}(x_{1},x_{2},\dots,x_{k}):=h_{1}(g_{x}x_{1})h_{2}(g_{x}x_{2})\cdots h_{k}(g_{x}x_{k})$. We then have
@@
In particular, this proves part 1. To prove part 2, suppose that condition 1 holds. Then $\nu_{x}$ is the Haar measure on $Y_{x}^{k}$, where $Y_{x}=\overline{\{g_{x}^{-1}a^{n}g_{x}\Gamma:n\in\mathbb{Z}\}}$. The Haar measure $\mu_{Y_{x}}$ on $Y_{x}$ can be described by $\int_{Y_{x}}h~{}\mathrm{d}\mu_{Y_{x}}=\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}f(g_{x}^{-1}a^{n}g_{x}\Gamma)$, and hence for each $i=1,\dots,k$,
@@
This together with (5.3) proves part 2.
To prove part 3 it is enough to show that for any continuous non-negative function $h\in C(X)$ with $\int_{X}h~{}\mathrm{d}\mu>0$ we have
@@ 5.4.   Removing the rounding function
Next, let us reduce 5.6 to a result that no longer involves the bracket function $[.]\colon\mathbb{R}\to\mathbb{Z}$.
@@
Let $G$ be a connected and simply connected nilpotent Lie group, let $\Gamma\subset G$ be a uniform and discrete subgroup, let $b\in G$ and $X=G/\Gamma$. Let $\mathcal{H}$ be a Hardy field and let $W\in\mathcal{H}$ satisfy $1\prec W(t)\ll t$. For any finite collection of functions $f_{1},\ldots,f_{k}\in\mathcal{H}$ of polynomial growth satisfying property 1 and property (5.1) there exists a Borel probability measure $\nu$ on $X^{k}$ such that the sequence $x_{n}=\big{(}b^{f_{1}(n)},\dots,b^{f_{k}(n)}\big{)}\Gamma^{k}$ is uniformly distributed with respect to $\nu$ and $W$-averages.
Moreover, if condition 2 of A is satisfied then the point $1_{X^{k}}=1_{G^{k}}\Gamma^{k}$ belongs to the support of $\nu$, and if condition 1 of A is satisfied then $\nu$ is the Haar measure on the subnilmanifold $Y^{k}$ where $Y\subset X$ is defined by $Y=\overline{\{b^{t}\Gamma:t\in\mathbb{R}\}}$.
Proof of 5.6 assuming 5.7.
@@
From 5.7 it follows that the sequence $x_{n}:=(b^{f_{1}(n)},b^{f_{2}(n)},\dots,b^{f_{k}(n)})\tilde{\Gamma}^{k}$ is uniformly distributed with respect to $\tilde{\nu}$ and $W$-averages, where $\tilde{\nu}$ is some Borel probability measure on $\tilde{X}^{k}$. We now consider two cases separately.
@@
    If condition 2 of A is satisfied then 5.7 also implies that the point $1_{\tilde{X}^{k}}$ is in the support of $\tilde{\nu}$. Since $\rho$ is continuous at 00, the map $\pi$ is continuous at $1_{\tilde{X}^{k}}$ and, because $\nu=\pi_{*}\tilde{\nu}$, it follows that the point $1_{X^{k}}=\pi(1_{\tilde{X}^{k}})$ belongs to the support of $\nu$.
    If condition 1 of A is satisfied then 5.7 also implies that $\tilde{\nu}$ is the Haar measure on the subnilmanifold $\tilde{Y}^{k}$ where $\tilde{Y}\subset\tilde{X}$ is defined by $\tilde{Y}=\overline{\{b^{t}\Gamma:t\in\mathbb{R}\}}$. Since $Y:=\{a^{n}\Gamma:n\in\mathbb{Z}\}$ satisfies $Y^{k}=\pi(\tilde{Y}^{k})$ it follows that $\nu=\pi_{*}\tilde{\nu}$ is the Haar measure on $Y^{k}$.
@@
    where $\|x\|_{\mathbb{T}}=|x-[x]|$ is the distance to the closest integer. In view of 2.5, there exists a polynomial $p_{i}\in\mathbb{Q}[x]$ such that $|f_{i}(n)-p_{i}(n)|\to 0$ as $n\to\infty$. In view of property (5.1), for each $i\leqslant k_{0}$ we in fact have $p_{i}(x)-p_{i}(0)\in\mathbb{Z}[t]$. This implies in particular that Case II is incompatible with condition 1 of A.
@@
    Let $y_{n}=(a^{f_{1}(n)},1/2,\dots,a^{f_{k_{0}}(n)},1/2,b^{f_{k_{0}+1}(n)},\dots,b^{f_{k}(n)})\tilde{\Gamma}^{k}\in D\subset\tilde{X}^{k}$ and note that $d(x_{n},y_{n})\to 0$ as $n\to\infty$ outside of a set of zero density, where $d$ is any compatible metric on $\tilde{X}^{k}$. It follows that $(y_{n})_{n\in\mathbb{N}}$ is also uniformly distributed with respect to $\tilde{\nu}$ and $W$-averages. Unfortunately, it is not necessarily the case that $\pi(y_{n})$ is getting close to $((a^{[f_{1}(n)]},\dots,a^{[f_{k}(n)]})\Gamma^{k})$ (see 5.8 below).
    Since for every $i\leqslant k_{0}$, the function $x\mapsto f_{i}(x)-p_{i}(x)$ is in $\mathcal{H}$, it eventually stops changing sign. Recalling that $p_{i}(n)-1/2$ is always an integer and $\rho$ has only jump discontinuities at $1/2+\mathbb{Z}$, we deduce that $\rho(f_{i}(n))=\rho(f_{i}(n)-p_{i}(n)+1/2)$ converges as $n\to\infty$ (to either $1/2$ or $-1/2$). Let
@@
    Since $[f_{i}(n)]=f_{i}(n)-\rho(f_{i}(n))$, in view of (5.5) after unraveling the definitions we have that
@@
    Finally, if $f_{1},\dots,f_{k}$ satisfy condition 2 of A, then arguing as in Case I we see that the point $1_{G^{k}}\Gamma^{k}=1_{X^{k}}$ belongs to the support of $\nu$. Since we already saw above that, in Case II, condition 1 of A can not hold, this finishes the proof.
@@
[…]
5.5.   A Proof of 5.7
The purpose of this subsection is to give a proof of 5.7. For this proof, we need two of the main results from [Ric23].
@@
Let $G$ be a connected and simply connected nilpotent Lie group, $\Gamma$ a uniform and discrete subgroup of $G$, and $\mathcal{H}$ a Hardy field. Assume $W\in\mathcal{H}$ satisfies $1\prec W(t)\ll t$ and $f_{1},\ldots,f_{k}\in\mathcal{H}$ satisfy property 1. Let
@@
Given a connected nilmanifold $X=G/\Gamma$, the maximal factor torus of $X$ is the quotient $[G^{\circ},G^{\circ}]\backslash X$, where $G^{\circ}$ is the identity component of $G$. We will use $\vartheta\colon X\to[G^{\circ},G^{\circ}]\backslash X$ to denote the factor map from $X$ onto $[G^{\circ},G^{\circ}]\backslash X$. Note that the maximal factor torus is the torus of maximal dimension that is a factor of the nilmanifold $X$.
@@
Let $G$ be a simply connected nilpotent Lie group, $\Gamma$ a uniform and discrete subgroup of $G$, and assume the nilmanifold $X=G/\Gamma$ is connected. Let $\mathcal{H}$ be a Hardy field, $W$ a function in $\mathcal{H}$ satisfying $1\prec W(t)\ll t$, and $f_{1},\ldots,f_{k}\in\mathcal{H}$ functions in $\mathcal{H}$ satisfying property 1. Consider
@@
  3. 2.
    The sequence $(\vartheta(v(n)\Gamma))_{n\in\mathbb{N}}$ is uniformly distributed with respect to $W$-averages in the maximal factor torus $[G^{\circ},G^{\circ}]\backslash X$.
We actually need the following corollary of 5.10.
@@
  1. [label=(),ref=(),leftmargin=*]
  2. 1.
    $\{g_{1},\ldots,g_{m}\}$ satisfies Property 1;
  3. 2.
@@
According to 5.10 it suffices to show that the projection of the sequence
@@
onto the maximal factor torus $[G^{\circ},G^{\circ}]\backslash X$ is uniformly distributed with respect to $W$-averages there. Since $X=G/\Gamma$ is connected, we have that $G^{\circ}\Gamma=\Gamma$. This implies (see [Lei05, Subsection 2.6]) that there exist $e_{1},\ldots,e_{s}\in G^{\circ}$ and $q_{1},\ldots,q_{s}\in\mathbb{Z}[t]$ such that
@@
is uniformly distributed with respect to $W$-averages in the maximal factor torus $[G^{\circ},G^{\circ}]\backslash X$. Note that since $X$ is connected, the embedding $G^{\circ}\hookrightarrow G$ descents to an isomorphism between the finite dimensional torus $G^{\circ}/([G^{\circ},G^{\circ}]\Gamma)$ and the quotient $[G^{\circ},G^{\circ}]\backslash X$. The down-side of replacing $b_{1}^{p_{1}(n)}\cdot\ldots\cdot b_{l}^{p_{l}(n)}\Gamma$ with $e_{1}^{q_{1}(n)}\cdot\ldots\cdot e_{s}^{q_{s}(n)}\Gamma$ is that the $e_{i}$s and $a_{j}$s are not necessarily commuting. However, after projecting onto $[G^{\circ},G^{\circ}]\backslash X$, this doesn’t matter. On the flip side, the advantage of replacing $b_{1}^{p_{1}(n)}\cdot\ldots\cdot b_{l}^{p_{l}(n)}\Gamma$ with $e_{1}^{q_{1}(n)}\cdot\ldots\cdot e_{s}^{q_{s}(n)}\Gamma$ is that the action of the sequence (5.7) remains an action by translation when passed through the isomorphism that identifies $[G^{\circ},G^{\circ}]\backslash X$ with $G^{\circ}/([G^{\circ},G^{\circ}]\Gamma)$ because its generators $a_{1},\ldots,a_{k},e_{1},\ldots,e_{s}$ belong to $G^{\circ}$. Working within translations on $G^{\circ}/([G^{\circ},G^{\circ}]\Gamma)$ is convenient since it enables us to use uniform distribution results for finite dimensional tori from Section 2.
@@
is dense, together with properties 1 and 2, the Weyl criterion and 2.3. ∎
Proof of 5.7.
@@
It follows directly from 5.9 that there exist $q\in\mathbb{N}$ and connected sub-nilmanifolds $Y_{0},Y_{1},\ldots,Y_{q-1}$ of $X^{k}$ such that for every $r\in\{0,1,\ldots,q-1\}$ the sequence

### v3

@@ 5.2.  Dealing with rational polynomials
˜5.3 holds if we further assume that each function $f_{i}$ satisfies the condition
[…]
To reduce ˜5.3 to ˜5.4 we need the following lemma.
@@
Because $W\in\mathcal{H}$, $\lim_{t\to\infty}W(t)/t$ exists and, because $W(t)\ll t$, that limit is in $0,\infty)$. If it is nonzero, then $L:=\lim\frac{W(Rt+d)}{W(t)}=\lim\frac{W(Rt+d)}{Rt+d}\frac{t}{W(t)}\frac{Rt+d}{t}=R$, and if $\lim W(t)/t=0$ then $W(t)/t$ is eventually decreasing, so $L\leqslant R$. Since $W$ is eventually increasing (because $1\prec W(t)$) we have $L\geqslant 1$. Next it follows from L’Hôpital’s rule and the mean value theorem that $w(Rt+d)/w(t)\to\frac{L}{R}$.
@@
Observe that the polynomials $p_{i,d}\colon t\mapsto p_{i}(Rt+d)-p_{i}(d)$ have integer coefficients for any fixed $d\in\mathbb{N}$ and any $i\leqslant m$. Therefore, for each $d\in\mathbb{N}$ and $i\in\{1,\dots,k\}$ the function $f_{i,d}\colon t\mapsto f_{i}(Rt+d)$ satisfies condition (5.1). Parts (i) and (ii) of ˜5.3 now follow from ˜5.4 and ˜5.5.
Finally, we prove part (iii). Suppose that the functions $f_{1},\ldots,f_{k}$ satisfy condition (2) of Theorem A and let $q_{1},\dots,q_{\ell}\in\mathbb{Z}[t]$ be jointly intersective and such that $\mathrm{poly}(f_{1},\dots,f_{k})\subset\mathrm{span}^{*}(q_{1},\dots,q_{\ell})$. By the pigeonhole principle, for some $d\in\{0,\dots,R-1\}$ the polynomials $\tilde{q}_{i}:t\mapsto q_{i}(Rt+d)$, $i=1,\dots,\ell$ are jointly intersective. Therefore the functions $f_{1,d},\ldots,f_{k,d}$ also satisfy condition (2) of Theorem A. Then using ˜5.5 and ˜5.4 we have
@@ 5.3.  Reduction to a statement about uniform distribution
Thus far, we have reduced ˜5.1 to ˜5.4. Our next goal is to reduce ˜5.4 to the following result:
@@
Let $G$ be a connected and simply connected nilpotent Lie group, let $\Gamma\subset G$ be a uniform and discrete subgroup, let $X=G/\Gamma$, and let $a\in G$. Let $f_{1},\dots,f_{k}\in\mathcal{H}$ satisfying property (P) for some $W\in\mathcal{H}$ with $1\prec W(t)\ll t$ and property (5.1) from ˜5.4. Then there exists a Borel probability measure $\nu$ on $X^{k}$ such that the sequence
@@
Moreover, if condition (2) of Theorem A is satisfied then the point $1_{X^{k}}=1_{G^{k}}\Gamma^{k}$ belongs to the support of $\nu$, and if condition (1) of Theorem A is satisfied then $\nu$ is the Haar measure on the subnilmanifold $Y^{k}\subset X^{k}$, where $Y=\overline{\{a^{n}\Gamma:n\in\mathbb{Z}\}}$.
Proof of ˜5.4 assuming ˜5.6.
Let $G$ be a connected and simply connected nilpotent Lie group and let $\Gamma\subset G$ be a uniform and discrete subgroup such that $X=G/\Gamma$ and $T:X\to X$ is given by left multiplication by some element $a\in G$. Our goal is to show that:
  1. (i)
@@
  2. (ii)
    If $f_{1},\ldots,f_{k}$ satisfy condition (1) of Theorem A then for any $h_{1},\ldots,h_{k}\in C(X)$ and any $x\in X$
@@
  3. (iii)
    If $f_{1},\ldots,f_{k}$ satisfy condition (2) of Theorem A then for any $A\in\mathcal{B}$ with $\mu(A)>0$ we have
    $\lim_{N\to\infty}\frac{1}{W(N)}\sum_{n=1}^{N}w(n)\mu\Big(A\cap T^{-[f_{1}(n)]}A\cap\ldots\cap T^{-[f_{k}(n)]}A\Big)\,>\,0.$
@@
Now let $\nu_{x}$ be the measure on $X^{k}$ given by ˜5.6 with $g_{x}^{-1}ag_{x}$ in the role of $a$ and define $H_{x}(x_{1},x_{2},\dots,x_{k}):=h_{1}(g_{x}x_{1})h_{2}(g_{x}x_{2})\cdots h_{k}(g_{x}x_{k})$. We then have
@@
In particular, this proves part (i). To prove part (ii), suppose that condition (1) holds. Then $\nu_{x}$ is the Haar measure on $Y_{x}^{k}$, where $Y_{x}=\overline{\{g_{x}^{-1}a^{n}g_{x}\Gamma:n\in\mathbb{Z}\}}$. The Haar measure $\mu_{Y_{x}}$ on $Y_{x}$ can be described by $\int_{Y_{x}}h~\mathrm{d}\mu_{Y_{x}}=\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}f(g_{x}^{-1}a^{n}g_{x}\Gamma)$, and hence for each $i=1,\dots,k$,
@@
This together with (5.3) proves part (ii).
To prove part (iii) it is enough to show that for any continuous non-negative function $h\in C(X)$ with $\int_{X}h~\mathrm{d}\mu>0$ we have
@@ 5.4.  Removing the rounding function
[…]
@@
Let $G$ be a connected and simply connected nilpotent Lie group, let $\Gamma\subset G$ be a uniform and discrete subgroup, let $b\in G$ and $X=G/\Gamma$. Let $\mathcal{H}$ be a Hardy field and let $W\in\mathcal{H}$ satisfy $1\prec W(t)\ll t$. For any finite collection of functions $f_{1},\ldots,f_{k}\in\mathcal{H}$ of polynomial growth satisfying property (P) and property (5.1) there exists a Borel probability measure $\nu$ on $X^{k}$ such that the sequence $x_{n}=\big(b^{f_{1}(n)},\dots,b^{f_{k}(n)}\big)\Gamma^{k}$ is uniformly distributed with respect to $\nu$ and $W$-averages.
Moreover, if condition (2) of Theorem A is satisfied then the point $1_{X^{k}}=1_{G^{k}}\Gamma^{k}$ belongs to the support of $\nu$, and if condition (1) of Theorem A is satisfied then $\nu$ is the Haar measure on the subnilmanifold $Y^{k}$ where $Y\subset X$ is defined by $Y=\overline{\{b^{t}\Gamma:t\in\mathbb{R}\}}$.
Proof of ˜5.6 assuming ˜5.7.
@@
From ˜5.7 it follows that the sequence $x_{n}:=(b^{f_{1}(n)},b^{f_{2}(n)},\dots,b^{f_{k}(n)})\tilde{\Gamma}^{k}$ is uniformly distributed with respect to $\tilde{\nu}$ and $W$-averages, where $\tilde{\nu}$ is some Borel probability measure on $\tilde{X}^{k}$. We now consider two cases separately.
@@
    If condition (2) of Theorem A is satisfied then ˜5.7 also implies that the point $1_{\tilde{X}^{k}}$ is in the support of $\tilde{\nu}$. Since $\rho$ is continuous at 0, the map $\pi$ is continuous at $1_{\tilde{X}^{k}}$ and, because $\nu=\pi_{*}\tilde{\nu}$, it follows that the point $1_{X^{k}}=\pi(1_{\tilde{X}^{k}})$ belongs to the support of $\nu$.
    If condition (1) of Theorem A is satisfied then ˜5.7 also implies that $\tilde{\nu}$ is the Haar measure on the subnilmanifold $\tilde{Y}^{k}$ where $\tilde{Y}\subset\tilde{X}$ is defined by $\tilde{Y}=\overline{\{b^{t}\Gamma:t\in\mathbb{R}\}}$. Since $Y:=\{a^{n}\Gamma:n\in\mathbb{Z}\}$ satisfies $Y^{k}=\pi(\tilde{Y}^{k})$ it follows that $\nu=\pi_{*}\tilde{\nu}$ is the Haar measure on $Y^{k}$.
@@
    where $\|x\|_{\mathbb{T}}=|x-[x]|$ is the distance to the closest integer. In view of ˜2.5, there exists a polynomial $p_{i}\in\mathbb{Q}[x]$ such that $|f_{i}(n)-p_{i}(n)|\to 0$ as $n\to\infty$. In view of property (5.1), for each $i\leqslant k_{0}$ we in fact have $p_{i}(x)-p_{i}(0)\in\mathbb{Z}[t]$. This implies in particular that Case II is incompatible with condition (1) of Theorem A.
@@
    Let $y_{n}=(a^{f_{1}(n)},1/2,\dots,a^{f_{k_{0}}(n)},1/2,b^{f_{k_{0}+1}(n)},\dots,b^{f_{k}(n)})\tilde{\Gamma}^{k}\in D\subset\tilde{X}^{k}$ and note that $d(x_{n},y_{n})\to 0$ as $n\to\infty$ outside of a set of zero density, where $d$ is any compatible metric on $\tilde{X}^{k}$. It follows that $(y_{n})_{n\in\mathbb{N}}$ is also uniformly distributed with respect to $\tilde{\nu}$ and $W$-averages. Unfortunately, it is not necessarily the case that $\pi(y_{n})$ is getting close to $((a^{[f_{1}(n)]},\dots,a^{[f_{k}(n)]})\Gamma^{k})$ (see ˜5.8 below).
    Since for every $i\leqslant k_{0}$, the function $x\mapsto f_{i}(x)-p_{i}(x)$ is in $\mathcal{H}$, it eventually stops changing sign. Recalling that $p_{i}(n)-1/2$ is always an integer and $\rho$ has only jump discontinuities at $1/2+\mathbb{Z}$, we deduce that $\rho(f_{i}(n))=\rho(f_{i}(n)-p_{i}(n)+1/2)$ converges as $n\to\infty$ (to either $1/2$ or 1/2-1/2). Let
@@
    Since $f_{i}(n)]=f_{i}(n)-\rho(f_{i}(n))$, in view of (5.5) after unraveling the definitions we have that
@@
    Finally, if $f_{1},\dots,f_{k}$ satisfy condition (2) of Theorem A, then arguing as in Case I we see that the point $1_{G^{k}}\Gamma^{k}=1_{X^{k}}$ belongs to the support of $\nu$. Since we already saw above that, in Case II, condition (1) of Theorem A can not hold, this finishes the proof.
@@
Let $f_{1}(n)=n$, $f_{2}(n)=n-1/n$, $W(n)=n$, $\cdot$$\lfloor\cdot\rfloor$, $G/\Gamma=\mathbb{T}=\mathbb{R}/\mathbb{Z}$ and $\a\in\mathbb{R}\backslash\mathbb{Q} a$rbitrary. Then $\tilde{G}=\mathbb{R}^{2}$, $\tilde{\Gamma}=\mathbb{Z}^{2}$ and $\pi:\mathbb{T}^{4}\to\mathbb{T}^{2}$ is given by $\pi(x_{1},t_{1},x_{2},t_{2})=(x_{1}-a\{t_{1}\},x_{2}-a\{t_{2}\})$ and $(b^{f_{1}(n)},b^{f_{2}(n)})\mathbb{Z}^{4}$ equidistributes on $Y=\{(t,0,t,0):t\in\mathbb{T}\}$. In this case $\pi$ is discontinuous on $Y$, so it falls into Case II of the proof. Indeed $\pi_{*}\mu_{Y}$ is the Haar measure on the diagonal $\{(x,x):x\in\mathbb{T}\}\subset\mathbb{T}^{2}$, but the sequence $(\lfloor n\rfloor a,\lfloor n-1/n\rfloor a)$ is uniformly distributed on the set $\{(x,x-a):x\in\mathbb{T}\}\subset\mathbb{T}^{2}$.
$5.$  A Proof of ˜5.7
The purpose of this subsection is to give a proof of ˜5.7. For this proof, we need two of the main results from [Ric23].
@@
Let $G$ be a connected and simply connected nilpotent Lie group, $\Gamma$ a uniform and discrete subgroup of $G$, and $\mathcal{H}$ a Hardy field. Assume $W\in\mathcal{H}$ satisfies $1\prec W(t)\ll t$ and $f_{1},\ldots,f_{k}\in\mathcal{H}$ satisfy property (P). Let
@@
Given a connected nilmanifold $X=G/\Gamma$, the maximal factor torus of $X$ is the quotient $\X[G^{\circ},G^{\circ}]\backslash$ X, where $G^{\circ}$ is the identity component of $G$. We will use $\X\vartheta\colon X\to[G^{\circ},G^{\circ}]\backslash$ X to denote the factor map from $X$ onto $\X[G^{\circ},G^{\circ}]\backslash$ X. Note that the maximal factor torus is the torus of maximal dimension that is a factor of the nilmanifold $X$.
@@
Let $G$ be a simply connected nilpotent Lie group, $\Gamma$ a uniform and discrete subgroup of $G$, and assume the nilmanifold $X=G/\Gamma$ is connected. Let $\mathcal{H}$ be a Hardy field, $W$ a function in $\mathcal{H}$ satisfying $1\prec W(t)\ll t$, and $f_{1},\ldots,f_{k}\in\mathcal{H}$ functions in $\mathcal{H}$ satisfying property (P). Consider
@@
  2. (ii)
    The sequence $(\vartheta(v(n)\Gamma))_{n\in\mathbb{N}}$ is uniformly distributed with respect to $W$-averages in the maximal factor torus $\X[G^{\circ},G^{\circ}]\backslash$ X.
We actually need the following corollary of ˜5.10.
@@
  1. (A)
    $\{g_{1},\ldots,g_{m}\}$ satisfies Property (P);
  2. (B)
@@
According to ˜5.10 it suffices to show that the projection of the sequence
@@
onto the maximal factor torus $\X[G^{\circ},G^{\circ}]\backslash$ X is uniformly distributed with respect to $W$-averages there. Since $X=G/\Gamma$ is connected, we have that $G^{\circ}\Gamma=\Gamma$. This implies (see [Lei05, Subsection 2.6]) that there exist $e_{1},\ldots,e_{s}\in G^{\circ}$ and $q_{1},\ldots,q_{s}\in\mathbb{Z}[t]$ such that
@@
is uniformly distributed with respect to $W$-averages in the maximal factor torus $\X[G^{\circ},G^{\circ}]\backslash$ X. Note that since $X$ is connected, the embedding $G^{\circ}\hookrightarrow G$ descents to an isomorphism between the finite dimensional torus $G^{\circ}/([G^{\circ},G^{\circ}]\Gamma)$ and the quotient $\X[G^{\circ},G^{\circ}]\backslash$ X. The down-side of replacing $b_{1}^{p_{1}(n)}\cdot\ldots\cdot b_{l}^{p_{l}(n)}\Gamma$ with $e_{1}^{q_{1}(n)}\cdot\ldots\cdot e_{s}^{q_{s}(n)}\Gamma$ is that the $e_{i}$s and $a_{j}$s are not necessarily commuting. However, after projecting onto $\X[G^{\circ},G^{\circ}]\backslash$ X, this doesn’t matter. On the flip side, the advantage of replacing $b_{1}^{p_{1}(n)}\cdot\ldots\cdot b_{l}^{p_{l}(n)}\Gamma$ with $e_{1}^{q_{1}(n)}\cdot\ldots\cdot e_{s}^{q_{s}(n)}\Gamma$ is that the action of the sequence (5.7) remains an action by translation when passed through the isomorphism that identifies $\X[G^{\circ},G^{\circ}]\backslash$ X with $G^{\circ}/([G^{\circ},G^{\circ}]\Gamma)$ because its generators $a_{1},\ldots,a_{k},e_{1},\ldots,e_{s}$ belong to $G^{\circ}$. Working within translations on $G^{\circ}/([G^{\circ},G^{\circ}]\Gamma)$ is convenient since it enables us to use uniform distribution results for finite dimensional tori from Section˜2.
@@
is dense, together with properties (A) and (B), the Weyl criterion and ˜2.3. ∎
Proof of ˜5.7.
@@
It follows directly from ˜5.9 that there exist $q\in\mathbb{N}$ and connected sub-nilmanifolds $Y_{0},Y_{1},\ldots,Y_{q-1}$ of $X^{k}$ such that for every $r\in\{0,1,\ldots,q-1\}$ the sequence